OpenStudy (anonymous):

The distribution of the number of occurrences of the letter t on the pages of a book is found to be a normal distribution with a mean of 44 and a standard deviation of 18. If there are 500 pages in the book, which sentence most closely summarizes the data? -The letter t occurs less than 26 times on approximately 170 of these pages. -The letter t occurs less than 26 times on approximately 15 of these pages. -The letter t occurs more than 26 times on approximately 420 of these pages. -The letter t occurs more than 26 times on approximately 80 of these pages.

OpenStudy (anonymous):

Help.!!

OpenStudy (anonymous):

Compute the probability that the letter occurs more less than 26 times, given by \[P(X<26)\] where \(X\) denotes the random variable for the number of times the letter t shows up. Use the variable transformation \(Z=\dfrac{X-\mu}{\sigma}\) to convert to the standard normal distribution, where \(\mu\) is the mean and \(\sigma\) is the standard deviation. \[P(X<26)=P\left(\frac{X-44}{18}<\frac{26-44}{18}\right)=P(Z<-1)\] Refer to a table to determine the probability, or use the rule here: http://en.wikipedia.org/wiki/68%E2%80%9395%E2%80%9399.7_rule The probability \(P(Z<-1)\) is the proportion of the population that lies at least one standard deviation to the left of the mean: